OpenStudy (anonymous):

Hue is 56 in. tall. His friend is 42 in. tall. Hue’s shadow is 24 in. long. How long is his friend’s shadow at the same time?

OpenStudy (anonymous):

24/56=x/42 x=24*42/56 x=18in.